# Kolmogorov-Arnold Energy Models: Fast and Interpretable Generative Modeling

### v8

@@ A.4.3 Maximum likelihood learning gradient
A.4.4 Contrastive divergence learning gradient for log-prior
The learning gradient attributed to the prior can be derived into a contrastive divergence framework, which is more typical of energy-based modeling literature. The prior model is solely dependent on $\bm{f}$ in Eq. 4. We can rewrite Eq. 17 as:
$\nabla_{\bm{f}}\log\left(p_{\bm{f}}\left(\bm{z}\right)\right)$ $=\nabla_{\bm{f}}\sum_{q=1}^{2n_{z}+1}\sum_{p=1}^{n_{z}}\log\left(\pi_{q,p}(z_{q,p})\right)=\nabla_{\bm{f}}\sum_{q=1}^{2n_{z}+1}\sum_{p=1}^{n_{z}}f_{q,p}(z_{q,p})+\log\left(\pi_{0}(z_{q,p})\right)-\log\left(Z_{q,p}\right)$ (39)
where $Z_{q,p}=\int_{\mathcal{Z}}\exp(f_{q,p}(z_{p}))\,d\pi_{0}(z_{p})$ is the normalization constant. We can simplify the constant using Eq. 36:
$\mathbb{E}_{p_{\bm{f}}(\bm{z})}\Bigg[\nabla_{\bm{f}}\log p_{\bm{f}}\left(\bm{z}\right)\Bigg]$ $=\mathbb{E}_{p_{\bm{f}}(\bm{z})}\Bigg[\nabla_{\bm{f}}\sum_{q=1}^{2n_{z}+1}\sum_{p=1}^{n_{z}}f_{q,p}(z_{q,p})+\log\left(\pi_{0}(z_{q,p})\right)-\log\left(Z_{q,p}\right)\Bigg]=0$
$=\mathbb{E}_{p_{\bm{f}}(\bm{z})}\Bigg[\nabla_{\bm{f}}\sum_{q=1}^{2n_{z}+1}\sum_{p=1}^{n_{z}}f_{q,p}(z_{q,p})\Bigg]-\nabla_{\bm{f}}\sum_{q=1}^{2n_{z}+1}\sum_{p=1}^{n_{z}}\log\left(Z_{q,p}\right)=0$
$\therefore\nabla_{\bm{f}}\sum_{q=1}^{2n_{z}+1}\sum_{p=1}^{n_{z}}\log\left(Z_{q,p}\right)$ $=\mathbb{E}_{p_{\bm{f}}(\bm{z})}\Bigg[\nabla_{\bm{f}}\sum_{q=1}^{2n_{z}+1}\sum_{p=1}^{n_{z}}f_{q,p}(z_{q,p})\Bigg]$ (40)
@@ A.4.4 Contrastive divergence learning gradient for log-prior
$\nabla_{\bm{f}}\log p_{\bm{f}}\left(\bm{z}\right)$ $=\nabla_{\bm{f}}\sum_{q=1}^{2n_{z}+1}\sum_{p=1}^{n_{z}}f_{q,p}(z_{q,p})-\mathbb{E}_{p_{\bm{f}}(\bm{z})}\Bigg[\nabla_{\bm{f}}\sum_{q=1}^{2n_{z}+1}\sum_{p=1}^{n_{z}}f_{q,p}(z_{q,p})\Bigg],$ (41)
@@
$\mathbb{E}_{p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x})}\Bigg[\nabla_{\bm{f}}\left[\log p_{\bm{f}}\left(\bm{z}\right)\right]\Bigg]=\;$ $\mathbb{E}_{p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x})}\Bigg[\nabla_{\bm{f}}\sum_{q=1}^{2n_{z}+1}\sum_{p=1}^{n_{z}}f_{q,p}(z_{q,p})\Bigg]$
-\; $\mathbb{E}_{p_{\bm{f}}(\bm{z})}\Bigg[\nabla_{\bm{f}}\sum_{q=1}^{2n_{z}+1}\sum_{p=1}^{n_{z}}f_{q,p}(z_{q,p})\Bigg].$ (42)
A.4.5 Deriving the Thermodynamic Integral
@@
$p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)},t)=\frac{p_{\bm{\Phi}}\left(\bm{x}^{(b)}\mid\bm{z}\right)^{t}\,\,p_{\bm{f}}(\bm{z})}{\bm{Z}_{t}}$ (43)
The partition function $\bm{Z}_{t}$ is given by:
$\bm{Z}_{t}=\mathbb{E}_{p_{\bm{f}}(\bm{z})}\left[p_{\bm{\Phi}}\left(\bm{x}^{(b)}\mid\bm{z}\right)^{t}\right]$ (44)
@@
$\frac{\partial}{\partial t}\log(\bm{Z}_{t})$ $=\frac{1}{\bm{Z}_{t}}\frac{\partial}{\partial t}\bm{Z}_{t}=\frac{1}{\bm{Z}_{t}}\mathbb{E}_{p_{\bm{f}}(\bm{z})}\left[\frac{\partial}{\partial t}p_{\bm{\Phi}}\left(\bm{x}^{(b)}\mid\bm{z}\right)^{t}\right]$
$=\frac{1}{\bm{Z}_{t}}\mathbb{E}_{p_{\bm{f}}(\bm{z})}\left[\log p_{\bm{\Phi}}\left(\bm{x}^{(b)}\mid\bm{z}\right)\cdot p_{\bm{\Phi}}\left(\bm{x}^{(b)}\mid\bm{z}\right)^{t}\right]\$
$=\mathbb{E}_{p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)},t)}\left[\log p_{\bm{\Phi}}\left(\bm{x}^{(b)}\mid\bm{z}\right)\right]$ (45)
@@
$\int_{0}^{1}\frac{\partial}{\partial t}\log(\bm{Z}_{t})\,dt=\int_{0}^{1}\mathbb{E}_{p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)},t)}\left[\log p_{\bm{\Phi}}\left(\bm{x}^{(b)}\mid\bm{z}\right)\right]\,dt$ (46)
Using the fundamental theorem of calculus, (larson2008calculus):
$\int_{0}^{1}\frac{\partial}{\partial t}\log(\bm{Z}_{t})\,dt=\log(\bm{Z}_{t=1})-\log(\bm{Z}_{t=0})=\int_{0}^{1}\mathbb{E}_{p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)},t)}\left[\log p_{\bm{\Phi}}\left(\bm{x}^{(b)}\mid\bm{z}\right)\right]\,dt$ (47)
@@
$\log(\bm{Z}_{t=0})$ $=\log\Big(\mathbb{E}_{p_{\bm{f}}(\bm{z})}\left[p_{\bm{\Phi}}\left(\bm{x}^{(b)}\mid\bm{z}\right)^{0}\right]\Big)=\log(1)=0$
$\log(\bm{Z}_{t=1})$ $=\log\Big(\mathbb{E}_{p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)})}\left[\log p_{\bm{\Phi}}\left(\bm{x}^{(b)}\mid\bm{z}\right)\right]\Big)=\log\mathbb{P}(\bm{x}^{(b)}\mid\bm{f},\bm{\Phi})$ (48)
@@
$\log(\bm{Z}_{t=1})-\log(\bm{Z}_{t=0})=\log\mathbb{P}(\bm{x}^{(b)}\mid\bm{f},\bm{\Phi})=\int_{0}^{1}\mathbb{E}_{p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)},t)}\left[\log p_{\bm{\Phi}}\left(\bm{x}^{(b)}\mid\bm{z}\right)\right]\,dt$ (49)
A.4.6 Discretizing the Thermodynamic Integral
Here, $t_{k}$ denotes the tempering at the $k$-th index of the schedule, and $N_{\text{t}}$ is the number of temperatures. As derived by CALDERHEAD20094028, Eq. 28 then be evaluated using:
$\log\left(p_{\bm{f},\bm{\Phi}}(\bm{x}^{(b)})\right)=\frac{1}{2}\sum_{k=1}^{N_{t}}\Delta t_{k}(E_{k-1}+E_{k})+\frac{1}{2}\sum_{k=1}^{N_{t}}D_{\text{KL}}(\mathbb{P}_{t_{k-1}}||\mathbb{P}_{t_{k}})-D_{\text{KL}}(\mathbb{P}_{t_{k}}||\mathbb{P}_{t_{k-1}})$ (50)
@@ A.4.6 Discretizing the Thermodynamic Integral
$\Delta t_{k}$ $=t_{k}-t_{k-1},\quad\quad E_{k}=\mathbb{E}_{p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)},t_{k})}\left[\log\left(p_{\bm{\Phi}}\left(\bm{x}^{(b)}\mid\bm{z}\right)\right)\right],$
$D_{\text{KL}}(\mathbb{P}_{t_{k-1}}||\mathbb{P}_{t_{k}})$ $=\mathbb{E}_{p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)},t_{k-1})}\Bigg[\log\Bigg(\frac{p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)},t_{k-1})}{p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)},t_{k})}\Bigg)\Bigg]$ (51)
This reflects the trapezium rule for numerical integration, supported by bias correction to provide an estimate of the Riemann integral that is especially robust to error. The learning gradient derived from the discretized Thermodynamic Integral is unchanged from the MLE gradient in Eq. 16, as demonstrated under Sec. A.4.7.
A.4.7 Learning gradient for discretized Thermodynamic Integral and power posterior partition function
@@
$D_{\text{KL}}(\mathbb{P}_{t_{k-1}}\|\mathbb{P}_{t_{k}})$ $=\mathbb{E}_{p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)},t_{k-1})}\Big[\log p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)},t_{k-1})-\log p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)},t_{k})\Big]$
$\text{and}\quad D_{\text{KL}}(\mathbb{P}_{t_{k}}\|\mathbb{P}_{t_{k-1}})$ $=\mathbb{E}_{p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)},t_{k})}\Big[\log p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)},t_{k})-\log p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)},t_{k-1})\Big]$
$=-\mathbb{E}_{p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)},t_{k})}\Big[\log p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)},t_{k-1})-\log p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)},t_{k})\Big]$ (52)
@@
$\log p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)},t_{k-1})$ $-\log p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)},t_{k})=-\Delta t_{k}\cdot\log p_{\bm{\Phi}}(\bm{x}^{(b)}\mid\bm{z})+\log\Bigg(\frac{\bm{Z}_{t_{k}}}{\bm{Z}_{t_{k-1}}}\Bigg)$ (53)
@@
$\therefore D_{\text{KL}}(\mathbb{P}_{t_{k-1}}\|\mathbb{P}_{t_{k}})$ $=-\Delta t_{k}\cdot\mathbb{E}_{p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)},t_{k-1})}\Big[\log p_{\bm{\Phi}}(\bm{x}^{(b)}\mid\bm{z})\Big]+\log\Bigg(\frac{\bm{Z}_{t_{k}}}{\bm{Z}_{t_{k-1}}}\Bigg)$
$=-\Delta t_{k}\cdot E_{k-1}+\log\Bigg(\frac{\bm{Z}_{t_{k}}}{\bm{Z}_{t_{k-1}}}\Bigg)$ (54)
$\text{and}\quad D_{\text{KL}}(\mathbb{P}_{t_{k}}\|\mathbb{P}_{t_{k-1}})$ $=\Delta t_{k}\cdot\mathbb{E}_{p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)},t_{k})}\Big[\log p_{\bm{\Phi}}(\bm{x}^{(b)}\mid\bm{z})\Big]-\log\Bigg(\frac{\bm{Z}_{t_{k}}}{\bm{Z}_{t_{k-1}}}\Bigg)$
$=\Delta t_{k}\cdot E_{k}-\log\Bigg(\frac{\bm{Z}_{t_{k}}}{\bm{Z}_{t_{k-1}}}\Bigg)$ (55)
@@
$\frac{1}{2}\sum_{k=1}^{N_{t}}D_{\text{KL}}(\mathbb{P}_{t_{k-1}}\|\mathbb{P}_{t_{k}})$ $-D_{\text{KL}}(\mathbb{P}_{t_{k}}\|\mathbb{P}_{t_{k-1}})=-\frac{1}{2}\sum_{k=1}^{N_{t}}\Delta t_{k}(E_{k-1}+E_{k})-2\log\Bigg(\frac{\bm{Z}_{t_{k}}}{\bm{Z}_{t_{k-1}}}\Bigg)$
$\therefore\log\left(p_{\bm{f},\bm{\Phi}}(\bm{x}^{(b)})\right)$ $=\frac{1}{2}\sum_{k=1}^{N_{t}}\Delta t_{k}(E_{k-1}+E_{k})+\frac{1}{2}\sum_{k=1}^{N_{t}}D_{\text{KL}}(\mathbb{P}_{t_{k-1}}||\mathbb{P}_{t_{k}})-D_{\text{KL}}(\mathbb{P}_{t_{k}}||\mathbb{P}_{t_{k-1}})$
$=\frac{1}{2}\sum_{k=1}^{N_{t}}\Delta t_{k}(E_{k-1}+E_{k})-\Bigg[\frac{1}{2}\sum_{k=1}^{N_{t}}\Delta t_{k}(E_{k-1}+E_{k})-2\log\Bigg(\frac{\bm{Z}_{t_{k}}}{\bm{Z}_{t_{k-1}}}\Bigg)\Bigg]$
$=\sum_{k=1}^{N_{t}}\log\Bigg(\frac{\bm{Z}_{t_{k}}}{\bm{Z}_{t_{k-1}}}\Bigg)=\sum_{k=1}^{N_{t}}\log(\bm{Z}_{t_{k}})-\log(\bm{Z}_{t_{k-1}})$ (56)
Following a similar method to Sec. A.4.4 we can use Eq. 36 to express the learning gradient attributed to each normalization constant in a simpler manner:
$\mathbb{E}_{p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)},t_{k})}\Big[\nabla_{\bm{f},\bm{\Phi}}\log p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)},t_{k})\Big]$
$=\,$ $\mathbb{E}_{p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)},t_{k})}\Big[t_{k}\cdot\nabla_{\bm{\Phi}}\log p_{\bm{\Phi}}(\bm{x}^{(b)}\mid\bm{z})+\nabla_{\bm{f}}\log p_{\bm{f}}(\bm{z})\Big]-\nabla_{\bm{f},\bm{\Phi}}\log(\bm{Z}_{t_{k}})=0$
$\therefore$ $\nabla_{\bm{f},\bm{\Phi}}\log(\bm{Z}_{t_{k}})=\mathbb{E}_{p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)},t_{k})}\Big[t_{k}\cdot\nabla_{\bm{\Phi}}\log p_{\bm{\Phi}}(\bm{x}^{(b)}\mid\bm{z})+\nabla_{\bm{f}}\log p_{\bm{f}}(\bm{z})\Big]$ (57)
@@
$\nabla_{\bm{f},\bm{\Phi}}\log\left(p_{\bm{f},\bm{\Phi}}(\bm{x}^{(b)})\right)$ $=\sum_{k=1}^{N_{t}}\nabla_{\bm{f},\bm{\Phi}}\log(\bm{Z}_{t_{k}})-\nabla_{\bm{f},\bm{\Phi}}\log(\bm{Z}_{t_{k-1}})$
$=\nabla_{\bm{f},\bm{\Phi}}\log(\bm{Z}_{t_{N_{t}}})-\nabla_{\bm{f},\bm{\Phi}}\log(\bm{Z}_{t_{0}})$
$=\mathbb{E}_{p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)})}\Bigg[\nabla_{\bm{f}}\log p_{\bm{f}}\left(\bm{z}\right)+\nabla_{\bm{\Phi}}\log p_{\bm{\Phi}}\left(\bm{x}^{(b)}\mid\bm{z}\right)\Bigg]$

### v9

@@ E.0.3 Maximum likelihood learning gradient
E.0.4 Contrastive divergence learning gradient for univariate log-prior
The learning gradient attributed to the univariate KAEM prior in Eq. 14 can be derived into a contrastive divergence framework, which is more typical of energy-based modeling literature. The prior model is solely dependent on $f$ in Eq. 14. We can rewrite Eq. 15 as:
$\nabla_{f}\log\left(p_{f}\left(\bm{z}\right)\right)$ $=\nabla_{f}\sum_{q=1}^{2n_{z}+1}\sum_{p=1}^{n_{z}}\log\left(p_{q,p}(z)\right)=\nabla_{f}\sum_{q=1}^{2n_{z}+1}\sum_{p=1}^{n_{z}}f_{q,p}(z)+\log\left(\pi_{0}(z)\right)-\log Z_{q,p}$ (44)
where $z=\int_{\mathcal{Z}}\exp(f_{q,p}(z_{p}))\,d\pi_{0}(z_{p})$ is the normalization constant. We can simplify the constant using Eq. 41:
$\mathbb{E}_{p_{f}(\bm{z})}\Bigg[\nabla_{f}\log p_{f}\left(\bm{z}\right)\Bigg]$ $=\mathbb{E}_{p_{f}(\bm{z})}\Bigg[\nabla_{f}\sum_{q=1}^{2n_{z}+1}\sum_{p=1}^{n_{z}}f_{q,p}(z)+\log\left(\pi_{0}(z)\right)-\log Z_{q,p}\Bigg]=0$
$=\mathbb{E}_{p_{f}(\bm{z})}\Bigg[\nabla_{f}\sum_{q=1}^{2n_{z}+1}\sum_{p=1}^{n_{z}}f_{q,p}(z)\Bigg]-\nabla_{f}\sum_{q=1}^{2n_{z}+1}\sum_{p=1}^{n_{z}}\log Z_{q,p}=0$
$\therefore\nabla_{f}\sum_{q=1}^{2n_{z}+1}\sum_{p=1}^{n_{z}}\log Z_{q,p}$ $=\mathbb{E}_{p_{f}(\bm{z})}\Bigg[\nabla_{f}\sum_{q=1}^{2n_{z}+1}\sum_{p=1}^{n_{z}}f_{q,p}(z)\Bigg]$ (45)
@@ E.0.4 Contrastive divergence learning gradient for univariate log-prior
$\nabla_{f}\log p_{f}\left(\bm{z}\right)$ $=\nabla_{f}\sum_{q=1}^{2n_{z}+1}\sum_{p=1}^{n_{z}}f_{q,p}(z)-\mathbb{E}_{p_{f}(\bm{z})}\Bigg[\nabla_{f}\sum_{q=1}^{2n_{z}+1}\sum_{p=1}^{n_{z}}f_{q,p}(z)\Bigg],$ (46)
@@
$\mathbb{E}_{p_{f,\Phi}(\bm{z}\mid\bm{x})}\Bigg[\nabla_{f}\left[\log p_{f}\left(\bm{z}\right)\right]\Bigg]=\;$ $\mathbb{E}_{p_{f,\Phi}(\bm{z}\mid\bm{x})}\Bigg[\nabla_{f}\sum_{q=1}^{2n_{z}+1}\sum_{p=1}^{n_{z}}f_{q,p}(z)\Bigg]$
-\; $\mathbb{E}_{p_{f}(\bm{z})}\Bigg[\nabla_{f}\sum_{q=1}^{2n_{z}+1}\sum_{p=1}^{n_{z}}f_{q,p}(z)\Bigg].$ (47)
E.0.5 Deriving the Thermodynamic Integral
@@
$p_{f,\Phi}(\bm{z}\mid\bm{x},t)=\frac{p_{\Phi}\left(\bm{x}\mid\bm{z}\right)^{t}\,\,p_{f}(\bm{z})}{Z_{t}}$ (48)
The partition function $Z_{t}$ is given by:
$Z_{t}=\mathbb{E}_{p_{f}(\bm{z})}\left[p_{\Phi}\left(\bm{x}\mid\bm{z}\right)^{t}\right]$ (49)
@@
$\frac{\partial}{\partial t}\log(Z_{t})$ $=\frac{1}{Z_{t}}\frac{\partial}{\partial t}Z_{t}=\frac{1}{Z_{t}}\mathbb{E}_{p_{f}(\bm{z})}\left[\frac{\partial}{\partial t}p_{\Phi}\left(\bm{x}\mid\bm{z}\right)^{t}\right]$
$=\frac{1}{Z_{t}}\mathbb{E}_{p_{f}(\bm{z})}\left[\log p_{\Phi}\left(\bm{x}\mid\bm{z}\right)\cdot p_{\Phi}\left(\bm{x}\mid\bm{z}\right)^{t}\right]\$
$=\mathbb{E}_{p_{f,\Phi}(\bm{z}\mid\bm{x},t)}\left[\log p_{\Phi}\left(\bm{x}\mid\bm{z}\right)\right]$ (50)
@@
$\int_{0}^{1}\frac{\partial}{\partial t}\log(Z_{t})\,dt=\int_{0}^{1}\mathbb{E}_{p_{f,\Phi}(\bm{z}\mid\bm{x},t)}\left[\log p_{\Phi}\left(\bm{x}\mid\bm{z}\right)\right]\,dt$ (51)
Using the fundamental theorem of calculus, (Larson and Edwards, 2008):
$\int_{0}^{1}\frac{\partial}{\partial t}\log(Z_{t})\,dt=\log(Z_{t=1})-\log(Z_{t=0})=\int_{0}^{1}\mathbb{E}_{p_{f,\Phi}(\bm{z}\mid\bm{x},t)}\left[\log p_{\Phi}\left(\bm{x}\mid\bm{z}\right)\right]\,dt$ (52)
@@
$\log(Z_{t=0})$ $=\log\Big(\mathbb{E}_{p_{f}(\bm{z})}\left[p_{\Phi}\left(\bm{x}\mid\bm{z}\right)^{0}\right]\Big)=\log(1)=0$
$\log(Z_{t=1})$ $=\log\Big(\mathbb{E}_{p_{f,\Phi}(\bm{z}\mid\bm{x})}\left[\log p_{\Phi}\left(\bm{x}\mid\bm{z}\right)\right]\Big)=\log p_{f,\Phi}(\bm{x})$ (53)
@@
$\log(Z_{t=1})-\log(Z_{t=0})=\log p_{f,\Phi}(\bm{x})=\int_{0}^{1}\mathbb{E}_{p_{f,\Phi}(\bm{z}\mid\bm{x},t)}\left[\log p_{\Phi}\left(\bm{x}\mid\bm{z}\right)\right]\,dt$ (54)
E.0.6 Discretizing the Thermodynamic Integral
Here, $t_{k}$ denotes the tempering at the $k$-th index of the schedule, and $N_{\text{t}}$ is the number of temperatures. As derived by Calderhead and Girolami (2009), Eq. 21 then be evaluated using:
$\log\left(p_{f,\Phi}(\bm{x})\right)=\frac{1}{2}\sum_{k=1}^{N_{t}}\Delta t_{k}(E_{k-1}+E_{k})+\frac{1}{2}\sum_{k=1}^{N_{t}}D_{\text{KL}}(p_{t_{k-1}}||p_{t_{k}})-D_{\text{KL}}(p_{t_{k}}||p_{t_{k-1}})$ (55)
@@ E.0.6 Discretizing the Thermodynamic Integral
$\Delta t_{k}$ $=t_{k}-t_{k-1},\quad\quad E_{k}=\mathbb{E}_{p_{f,\Phi}(\bm{z}\mid\bm{x},t_{k})}\left[\log\left(p_{\Phi}\left(\bm{x}\mid\bm{z}\right)\right)\right],$
$D_{\text{KL}}(p_{t_{k-1}}||p_{t_{k}})$ $=\mathbb{E}_{p_{f,\Phi}(\bm{z}\mid\bm{x},t_{k-1})}\Bigg[\log\Bigg(\frac{p_{f,\Phi}(\bm{z}\mid\bm{x},t_{k-1})}{p_{f,\Phi}(\bm{z}\mid\bm{x},t_{k})}\Bigg)\Bigg]$ (56)
This reflects the trapezium rule for numerical integration, supported by bias correction to provide an estimate of the Riemann integral across $t$ that is especially robust to error. The learning gradient derived from the discretized Thermodynamic Integral is unchanged from the MLE gradient in Eq. 4, as demonstrated under Sec. E.0.7.
E.0.7 Learning gradient for discretized Thermodynamic Integral and power posterior partition function
@@
$D_{\text{KL}}(p_{t_{k-1}}\|p_{t_{k}})$ $=\mathbb{E}_{p_{f,\Phi}(\bm{z}\mid\bm{x},t_{k-1})}\Big[\log p_{f,\Phi}(\bm{z}\mid\bm{x},t_{k-1})-\log p_{f,\Phi}(\bm{z}\mid\bm{x},t_{k})\Big]$
$\text{and}\quad D_{\text{KL}}(p_{t_{k}}\|p_{t_{k-1}})$ $=\mathbb{E}_{p_{f,\Phi}(\bm{z}\mid\bm{x},t_{k})}\Big[\log p_{f,\Phi}(\bm{z}\mid\bm{x},t_{k})-\log p_{f,\Phi}(\bm{z}\mid\bm{x},t_{k-1})\Big]$
$=-\mathbb{E}_{p_{f,\Phi}(\bm{z}\mid\bm{x},t_{k})}\Big[\log p_{f,\Phi}(\bm{z}\mid\bm{x},t_{k-1})-\log p_{f,\Phi}(\bm{z}\mid\bm{x},t_{k})\Big]$ (57)
@@
$\log p_{f,\Phi}(\bm{z}\mid\bm{x},t_{k-1})$ $-\log p_{f,\Phi}(\bm{z}\mid\bm{x},t_{k})=-\Delta t_{k}\cdot\log p_{\Phi}(\bm{x}\mid\bm{z})+\log\Bigg(\frac{Z_{t_{k}}}{Z_{t_{k-1}}}\Bigg)$ (58)
@@
$\therefore D_{\text{KL}}(p_{t_{k-1}}\|p_{t_{k}})$ $=-\Delta t_{k}\cdot\mathbb{E}_{p_{f,\Phi}(\bm{z}\mid\bm{x},t_{k-1})}\Big[\log p_{\Phi}(\bm{x}\mid\bm{z})\Big]+\log\Bigg(\frac{Z_{t_{k}}}{Z_{t_{k-1}}}\Bigg)$
$=-\Delta t_{k}\cdot E_{k-1}+\log\Bigg(\frac{Z_{t_{k}}}{Z_{t_{k-1}}}\Bigg)$ (59)
$\text{and}\quad D_{\text{KL}}(p_{t_{k}}\|p_{t_{k-1}})$ $=\Delta t_{k}\cdot\mathbb{E}_{p_{f,\Phi}(\bm{z}\mid\bm{x},t_{k})}\Big[\log p_{\Phi}(\bm{x}\mid\bm{z})\Big]-\log\Bigg(\frac{Z_{t_{k}}}{Z_{t_{k-1}}}\Bigg)$
$=\Delta t_{k}\cdot E_{k}-\log\Bigg(\frac{Z_{t_{k}}}{Z_{t_{k-1}}}\Bigg)$ (60)
@@
$\frac{1}{2}\sum_{k=1}^{N_{t}}D_{\text{KL}}(p_{t_{k-1}}\|p_{t_{k}})$ $-D_{\text{KL}}(p_{t_{k}}\|p_{t_{k-1}})=-\frac{1}{2}\sum_{k=1}^{N_{t}}\Delta t_{k}(E_{k-1}+E_{k})-2\log\Bigg(\frac{Z_{t_{k}}}{Z_{t_{k-1}}}\Bigg)$
$\therefore\log\left(p_{f,\Phi}(\bm{x})\right)$ $=\frac{1}{2}\sum_{k=1}^{N_{t}}\Delta t_{k}(E_{k-1}+E_{k})+\frac{1}{2}\sum_{k=1}^{N_{t}}D_{\text{KL}}(p_{t_{k-1}}||p_{t_{k}})-D_{\text{KL}}(p_{t_{k}}||p_{t_{k-1}})$
$=\frac{1}{2}\sum_{k=1}^{N_{t}}\Delta t_{k}(E_{k-1}+E_{k})-\Bigg[\frac{1}{2}\sum_{k=1}^{N_{t}}\Delta t_{k}(E_{k-1}+E_{k})-2\log\Bigg(\frac{Z_{t_{k}}}{Z_{t_{k-1}}}\Bigg)\Bigg]$
$=\sum_{k=1}^{N_{t}}\log\Bigg(\frac{Z_{t_{k}}}{Z_{t_{k-1}}}\Bigg)=\sum_{k=1}^{N_{t}}\log(Z_{t_{k}})-\log(Z_{t_{k-1}})$ (61)
Following a similar method to Sec. E.0.4 we can use Eq. 41 to express the learning gradient attributed to each normalization constant in a simpler manner:
$\mathbb{E}_{p_{f,\Phi}(\bm{z}\mid\bm{x},t_{k})}\Big[\nabla_{f,\Phi}\log p_{f,\Phi}(\bm{z}\mid\bm{x},t_{k})\Big]$
$=\,$ $\mathbb{E}_{p_{f,\Phi}(\bm{z}\mid\bm{x},t_{k})}\Big[t_{k}\cdot\nabla_{\Phi}\log p_{\Phi}(\bm{x}\mid\bm{z})+\nabla_{f}\log p_{f}(\bm{z})\Big]-\nabla_{f,\Phi}\log(Z_{t_{k}})=0$
$\therefore$ $\nabla_{f,\Phi}\log(Z_{t_{k}})=\mathbb{E}_{p_{f,\Phi}(\bm{z}\mid\bm{x},t_{k})}\Big[t_{k}\cdot\nabla_{\Phi}\log p_{\Phi}(\bm{x}\mid\bm{z})+\nabla_{f}\log p_{f}(\bm{z})\Big]$ (62)
@@
$\nabla_{f,\Phi}\log\left(p_{f,\Phi}(\bm{x})\right)$ $=\sum_{k=1}^{N_{t}}\nabla_{f,\Phi}\log(Z_{t_{k}})-\nabla_{f,\Phi}\log(Z_{t_{k-1}})$
$=\nabla_{f,\Phi}\log(Z_{t_{N_{t}}})-\nabla_{f,\Phi}\log(Z_{t_{0}})$
$=\mathbb{E}_{p_{f,\Phi}(\bm{z}\mid\bm{x})}\Bigg[\nabla_{f}\log p_{f}\left(\bm{z}\right)+\nabla_{\Phi}\log p_{\Phi}\left(\bm{x}\mid\bm{z}\right)\Bigg]$